OpenStudy (anonymous):

Choose the number sentence that shows the identity property of addition.Select one of the options below as your answer: A. 34 + 34 = 68 B. 1 + 34 = 35 C. 0 + 34 = 34

OpenStudy (anonymous):

C. 0 + 34 = 34

OpenStudy (anonymous):

how?

Directrix (directrix):

Additive Identity Property: The sum of any number and zero is the original number. For example 5 + 0 = 5. http://www.aaastudy.com/add74ax1.htm

OpenStudy (anonymous):

so what about the identity property of multiplication

Directrix (directrix):

Identity property of multiplication states that the product of 1 and any number or variable is the number or variable itself. For example, 4 × 1 = 4, - 11 × 1 = - 11, y × 1 = y are few examples illustrating the identity property of multiplication. http://www.northstarmath.com/sitemap/identitypropertiesofadditionandmultiplication.html

OpenStudy (anonymous):

thank you so much, really appreciate it